OpenStudy (anonymous):

Can you please solve this equation and show the work? 25^1-3x = 125^x-1

OpenStudy (anonymous):

\[25^{1-3x} = 125^{x-1}\] is the right equation...

OpenStudy (anonymous):

@Zale101 @Data_LG2

OpenStudy (anonymous):

HI sorry for the late response.. first thing you have to do is to make the base of terms equal the common base for 25 and 125 would be 5 so you'll have: \(\sf 25^{1-3x} = 125^{x-1}\\ 5^{2(1-3x)}=5^{3(x-1)}\) since both side have the same base already, you can easily solve x in the exponents algebraically: \(\sf 2(1-3x)=3(x-1)\) can you do it from here?

OpenStudy (anonymous):

X is then equal to 5/9 right?

OpenStudy (anonymous):

\(\sf 2(1-3x)=3(x-1)\\2-6x=3x-3\\3x+6x=2+3\\9x=5\\x=5/9\) yes, looks like you're right :)

OpenStudy (anonymous):

Thanks!